OpenStudy (anonymous):

Factor: x2 - 8x + 16 (x - 4)(x + 4) (x - 8)(x + 8) (x - 8)2 (x - 4)2

OpenStudy (anonymous):

\[x^2-2xy+y^2=(x-y)^2\] \[(x)^2-2(4)(x)+(4)^2=(x-4)^2\]